OpenStudy (kittykr98):

use the function y+2=-(x-3)^2 to answer the following. A) show or explain how you would find the vertex. B) show or explain how to find the axis of symmetry. C) show or explain how you would find the x-intercepts. D) show or explain how you would find the y-intercept. E) Graph the function. I know that i first need to convert the function to vertex form y=(x-h)^2+k by subtracting 2 from both sides. which makes it y=-(x-3)^2 -2

OpenStudy (kittykr98):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`I know that i first need to convert the function to vertex form y=(x-h)^2+k by subtracting 2 from both sides. which makes it y=-(x-3)^2 -2` a great start so far

OpenStudy (kittykr98):

And i know that the vertex is (3,-2) but i dont know how to show work for how i got that answer

jimthompson5910 (jim_thompson5910):

there's not much work to show really. You're comparing that equation you got with the template `y = a(x-h)^2 + k` `(h,k)` is the vertex

OpenStudy (kittykr98):

so to answer the question i write, "i first need to convert the function to vertex form y=(x-h)^2+k by subtracting 2 from both sides. which makes it y=-(x-3)^2 -2 we know that (h,k) is the vertex, so the vertex is (3,-2)" ?

jimthompson5910 (jim_thompson5910):

I would have "we know that (h,k) is the vertex of y = a(x-h)^2 + k" or something like that

OpenStudy (kittykr98):

we know that (h,k) is the vertex of y=a(x-h)^2+k so in the case of y=-(x-3)^2-2, the vertex is (3,-2) better?

jimthompson5910 (jim_thompson5910):

yeah that looks good

jimthompson5910 (jim_thompson5910):

`B) show or explain how to find the axis of symmetry.` the axis of symmetry of `y=a(x-h)^2+k` is `x = h`. It is the vertical line through vertex.

OpenStudy (kittykr98):

so the axis of symmetry for y=-(x-3)^2 -2 is x=3

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (kittykr98):

or -3?

jimthompson5910 (jim_thompson5910):

x = 3 since h = 3

OpenStudy (kittykr98):

x intercepts

jimthompson5910 (jim_thompson5910):

To find the x-intercepts, plug in y = 0 and solve for x

OpenStudy (kittykr98):

wait hold on, x=3 because h=3 ? what?

jimthompson5910 (jim_thompson5910):

the vertex in general is (h,k) in this case, the vertex is (3,-2) notice how the h and 3 come first in the ordered pairs, so h = 3

jimthompson5910 (jim_thompson5910):

in general, the axis of symmetry is x = h in this problem, the axis of symmetry is x = 3

OpenStudy (kittykr98):

so i write "the AOS of y=a(x-h)^2+k is x=h. it is the vertical line through the vertex. so the AOS for y=-(x-3)^2-2 is 3, making x=3" is that what i should have written?

jimthompson5910 (jim_thompson5910):

yeah something like that

OpenStudy (kittykr98):

Ok. Then back to x intercepts. make y=0

jimthompson5910 (jim_thompson5910):

then solve for x

OpenStudy (kittykr98):